xy=8
y=1/2X-2
substitute or eliminate to isolate one variable

Answer by Boreal(15235)

xy=8
y=1/(2X-2)
x/(2x-2)=8; note, I am assuming this is the right formula given above and not (1/2)x-2
multiply by 2x-2
x=16x-16
-15x=-16
x=(16/15)
y=1/[(32-30/15]=1/(2/15)
=15/2
-----------------------
(16/15)(15/2)=8